OpenStudy (anonymous):

From 9 names on a ballot a committee of 4 will be elected to attened a political national convention. How many differnet committees are possible

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

36

OpenStudy (anonymous):

My answer says 126

OpenStudy (raden):

9C4

OpenStudy (anonymous):

What's the c

OpenStudy (raden):

COMBINATION remember it the fomula : nCr = n!/{(n-r)!r!}

OpenStudy (anonymous):

I get that but it doesn't match with the answer of 126 that I have to have

OpenStudy (raden):

9C4 = 9!/(9-4)!4! = 9!/5!4! = 9*8*7*6*5!/(5!*4*3*2*1) = 9*8*7*6 / 4*3*2*1 = 126